Succeeded in handing off the servo from the green to the red.
This time we found that the fluctuation in the IR signals became lesser as the gain of the ALS servo increased.
Therefore I increased the UGF from 40 Hz to 180 Hz to have less noise in the IR PDH signal.
Here is a preliminary plot for today's noise spectra.
The blue curve is the ALS in-loop spectrum, that corresponds to the beat fluctuation.
The red curve is an out-of-loop spectrum taken by measuring the IR PDH signal.
Since the UGF is at about 180 Hz the rms is integrated from 200 Hz.
The residual displacement noise in the IR PDH signal is now 1.2 kHz in rms.
I am going to analyze this residual noise by comparing with the differential noise that I took yesterday (see the last entry ).